OpenStudy (anonymous):

the linear equation 3x-y=-7 is written in standard form. identify the value of a. a)3 b)-1 c) -7 d)7

OpenStudy (anonymous):

the standard form is y = ax + b ... or ax + b = y. so we shuffle the y to one side and the 7 to the other 3x + 7 = y. giving (a=3) and (b=7)

OpenStudy (anonymous):

the answer for a is not 7 but 3. so is the answer 3 or 7

OpenStudy (anonymous):

3

OpenStudy (anonymous):

You're confusing the "a), b), c), and d)" of the multiple choice question with the "a (coefficient of x) and b (constant) in the standard form of the equation. Make sure you can understand the different steps of the problem - the math problem and the test answering problem. (Quadratic equations will be more fun - they will have a, b, and c constants in the equation !)

OpenStudy (anonymous):

The standard form is ax + by + c = 0 Here the equation is 3x + (-1)y + 7 = 0, so a=3 So solution is a) 3